OpenStudy (anonymous):

In the rhombus m<1=8y-6. Find the value of y

jhonyy9 (jhonyy9):

m<1 =8y-6 --- 1=8y-6 --- 7=8y --- y = 7/8 --- 7/8 < 1

OpenStudy (anonymous):

m<8y=7 m<1, y=7/8